OpenStudy (anonymous):

a store mixes red fescue at 411 per pound and ryegrass worth $14 per pound. The mixture is to sell for $12 perpound. Find how much each should be used to make a 408 pound mixture

OpenStudy (anonymous):

*$11 per pound

OpenStudy (dangerousjesse):

Okay, I read the question wrong haha.

OpenStudy (anonymous):

I got it a+b=408 11a+14b=12 408 11a+14b=4896 -11a-11b=4488 3b=408 b=136 a+136=408 a=408-136 a=272

OpenStudy (dangerousjesse):

\[408\div2=x\]now solve for x

OpenStudy (dangerousjesse):

Ah, nice job.

OpenStudy (anonymous):

Thank you. this was a hard question..